OpenStudy (anonymous):

how many places will the graph y = x^2-6x+6 intersect the axis?

OpenStudy (anonymous):

two

OpenStudy (anonymous):

so how did you get that?

OpenStudy (anonymous):

putting y=0 x^2-6x+6=0 so\[x=\frac{6\pm \sqrt{6^2-24}}{2}\]

OpenStudy (anonymous):

no the ans is not (x-3)(x-3)=0 just check within the radical i have 6^2-24=12 so x=3+-sqrt(3)

OpenStudy (anonymous):

yes dipank you are right - i had a mental block there - thats why i deleted the reply lol